OpenStudy (anonymous):

Find the equation of the line tangent to the curve xy^4+xy=780 at the point (3,4). First, you want to find an expression for dy/dx in terms of x and y using the implicit differentiation: i got (-y^4-y)/(x (4 y^3+1)) Then, the equation of the tangent line could be written as y=?

hartnn (hartnn):

u put the point (3,4) in dy/dx x=3, y=4

hartnn (hartnn):

then dy/dx will give you slope of tangent

OpenStudy (anonymous):

(-4^4-4)/(3 (4 4^3+1))=-260/771

OpenStudy (anonymous):

(y-4)=(-260/771)(x-3)=y-4=mx+260/257

hartnn (hartnn):

(y-4)=(-260/771)(x-3) is correct

OpenStudy (anonymous):

y=(-260/771)x+1288/257

OpenStudy (anonymous):

sweet thanks

hartnn (hartnn):

welcome ^_^